### v2

# Holographic Boundary Conformal Field Theory within Horndeski Gravity

Abstract
We investigate entanglement islands and the Page curve in the framework of Horndeski gravity on a Karch-Randall braneworld background. In particular, treating the holographic boundary conformal field theory analytically we find that the Horndeski parameters significantly alter the behavior of the Page curve compared to standard general relativity, a feature caused by the nontrivial geometry induced by the Horndeski scalar field. Interestingly enough, the geometry far from the AdS limit plays a more significant role compared to previous studies. This suggests that Horndeski gravity introduces important modifications to the distribution of quantum information in the holographic model. Finally, we claim that holographic consistency can be used reversely to impose constraints on Horndeski gravity itself, providing a new tool for probing the validity of modified gravity theories.
1 Introduction
The black hole information loss paradox has been a central issue in theoretical physics since its introduction, gaining particular prominence following the insights provided by Page curve. This curve suggests a potential violation of unitarity when the fine-grained entropy of Hawking radiation from an evaporating black hole surpasses the entropy of the black hole itself Hawking:1975vcx ; Hawking:1976ra . In response, a variety of alternative theoretical models have been proposed to address this paradox more effectively. Among these, the concept of holographic complexity has emerged as a significant approach Susskind:2014rva ; Brown:2015bva ; Lloyd:2000cry ; Brown:2015lvg ; Susskind:2018fmx ; Brown:2018bms ; Brown:2017jil ; Brown:2019whu ; Brown:2022rwi . This framework posits that black holes continue to emit information even after reaching thermal equilibrium.
Recent investigations have further explored these issues within the context of modified gravity theories, particularly Horndeski gravity Horndeski:1974wa ; Kobayashi:2019hrl ; Lu:2020iav ; Kase:2018aps ; Koyama:2013paa ; Bahamonde:2021dqn ; Petronikolou:2021shp , which is a significant subject knowing the theoretical and observational advantages of gravitational modifications CANTATA:2021asi ; Abdalla:2022yfr ; Capozziello:2011et ; Cai:2015emx . Studies such as those by Santos:2021orr ; Sokoliuk:2022llp ; Santos:2023flb ; Santos:2023mee have demonstrated that residual information can be extracted by post-thermal equilibrium through the Anti-de-Sitter/Boundary Conformal Field Theory (AdS/BCFT) correspondence. This correspondence underscores the critical role of black hole entropy (SBH) Santos:2024zoh in the black hole information paradox, offering a promising avenue for further research into the intricate relationship between black hole thermodynamics and holographic principle.
Hence, developments in holographic transport and gravity-related quantum effects present a comprehensive overview of the intricate connections between the Anti-de-Sitter/Conformal Field Theory (AdS/CFT) correspondence and modified gravity theories, particularly Horndeski gravity Maldacena:1997re ; Witten:1998qj ; Caceres:2023gfa ; Santos:2024cwf ; Takayanagi:2011zk ; Fujita:2011fp ; Kanda:2023zse ; dosSantos:2022scy . Moreover, the extension of AdS/CFT proposed by Takayanagi, which emphasizes boundary effects and their relation to entanglement entropy Takayanagi:2011zk ; Fujita:2011fp ; Ryu:2006bv ; DosSantos:2022exb ; Caceres:2017lbr ; Feng:2015oea ; Wald:1993nt ; Iyer:1994ys , opens promising avenues for exploring the interplay between quantum gravity and field theories Karch:2000ct ; DeWolfe:2001pq ; Bak:2003jk ; Clark:2004sb ; Cardy:2004hm ; Tonni:2010pv ; Azeyanagi:2007qj ; Setare:2008hm ; Saridakis:2017rdo ; Basilakos:2023seo . The discussion of Hawking-Page phase transitions and corrections to boundary entropy underscores the richness of the holographic framework in addressing complex phenomena. By describing gravity duals within this context enlightens the discussion on the relation between quantum information theory and gravitational physics, particularly in scenarios where traditional quantum gravity approaches may be inadequate.
We mention here that the entanglement entropy in the context of Horndeski gravity Ryu:2006bv ; DosSantos:2022exb aligns with the boundary entropy in two-dimensional BCFTs Santos:2021orr . In these frameworks, gravity duals for Einstein and Horndeski gravity are defined at the CFT boundary on the AdSd-dimensional manifold $\mathcal{M}$, which is asymptotically AdSd+1-dimensional. For visualization we refer to Fig. 1 to elucidate the topological aspects of the AdS space, where as we can see the involved boundaries provides a concrete foundation for the discussion, enhancing the understanding of these complex theoretical constructs.
Figure 1: Shcematic representation of AdS/BCFT correspondence. Here, $\partial\mathcal{N}=\mathcal{M}\cup\,Q$, where $Q$ is a $d$-dimensional manifold satisfying $\partial Q\cap\partial\mathcal{M}=\mathcal{P}$.
The recent discussions on doubly holographic descriptions and the island formula reveal the interplay between holography, entanglement, and gravity. The use of BCFTs to understand these concepts is a significant advancement, especially in the context of understanding how entanglement entropy behaves in gravitational settings Almheiri:2019hni ; Rozali:2019day ; Chen:2020uac ; Chen:2020hmv ; Deng:2020ent ; Suzuki:2022xwv ; Geng:2021iyq ; Geng:2024xpj ; Geng:2020fxl . The framework we are outlining, where a CFT coupled to semiclassical gravity provides a dual description of a (d+1)-dimensional theory, is a robust platform for exploring entanglement phenomena. Additionally, the Ryu-Takayanagi (RT) prescription is central to this exploration Ryu:2006ef ; Ryu:2006bv , enabling the calculation of holographic entanglement entropy in terms of geometric quantities in the gravitational dual Suzuki:2022xwv .
Our approach, deriving the island formula within Horndeski gravity and connecting it to work on holographic entanglement entropy in disjoint subsystems, offers insight to these entanglement measures. Hence, the treatment of subsystems $A$ (subregion $\mathcal{R}$) and its complement $\bar{\mathcal{R}}$, and their connection through the bifurcation surface, illustrates how the RT formula can be adapted in more complex gravitational settings Basu:2023jtf . The visualization using the Penrose diagram (see Fig. 2) to illustrate the relation between past and future horizons at the bifurcation point emphasizes the geometric intuition behind the island formula and its relevance in quantum gravity scenarios. Furthermore, this approach helps to clarify the island formula’s implications and underlines the importance of horizon dynamics in the context of entanglement and information loss. In particular, in this scenario we calculate the entropy of the subregion $\mathcal{R}$, which is connected internally through the bifurcation where the horizon at $t=0$ separates the past and future horizons (this is the point in the center of the Penrose diagram in Fig. 2).
Figure 2: Subsystems Penrose diagram (see text).
The exploration of gravitational duals in AdSd+1 through the framework of Karch-Randall (KR) branes presents a compelling approach to understanding the interplay between gravity on the brane and non-gravitational baths via holography Randall:1999vf ; Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp ; Almheiri:2019psf ; Penington:2019npb ; Engelhardt:2014gca ; Almheiri:2019hni ; Almheiri:2019psy . This connection to AdSd gravity, where the brane is coupled to a CFT, serves as a powerful tool for probing the dynamics of the system, particularly the entanglement entropy and information flow. Thus, the new input gained from Horndeski gravity, particularly the reduction of brane physics to entanglement entropy, is useful to understand the local physics on the brane Karch:2000ct . Traditional limits, such as those in disk AdS4-dS4, may not fully encapsulate the subtleties necessary to understand local phenomena comprehensively.
In the context of Santos:2023eqp , dividing mass excitations into two distinct sets provides a new perspective on the distribution and processing of quantum information between the CFT on the brane and the bulk dynamics. This separation has significant implications for understanding how information is stored and retrieved in these holographic models, especially in scenarios involving black holes and entropy dynamics.
The analysis of the interplay between a CFTd coupled to AdSd gravity on the brane, particularly in the context of transparent boundary conditions, offers information on the dynamics of such systems Basu:2023jtf ; Hartman:2013qma ; Geng:2024xpj ; Geng:2021mic ; Geng:2023qwm ; Geng:2021hlu ; Geng:2023zhq ; Geng:2022dua . These boundary conditions serve as a crucial interface between the brane dynamics and the non-gravitational bath, leading to significant physical phenomena, such as the effective mass acquisition of the graviton due to energy exchange Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp . The role of transparent boundary conditions in facilitating this mass acquisition is important, as highlighted in works like Geng:2021mic . This mechanism illustrates how the gravitational theory on the brane can influence bulk dynamics, thereby affecting the properties of gravitational excitations.
A significant observation is that these boundary conditions result in the non-conservation of the stress tensor within the BCFT defect Santos:2023mee . This non-conservation can be interpreted as a reflection of the dynamics of quantum information and entanglement in these settings, which is essential for understanding the implications of residual information within the Horndeski framework. Furthermore, linking boundary entropy to this residual information and its role in the growth of entanglement entropy reveals the relations present in these holographic scenarios Santos:2021orr ; Santos:2023mee . Thus, investigating how these two classes of surfaces contribute to the growth of entanglement entropy offers information on the relation between boundary conditions, entanglement dynamics, and gravitational theories.
Recently, some works investigated the entanglement islands and the Page curve within the framework of Horndeski Geng:2021mic ; Geng:2023qwm ; Geng:2021hlu . Our goal to explore entanglement islands within a Horndeski gravitational framework, focusing on a single scalar field that induces symmetry breaking, aligns with studies suggesting that weakly broken symmetries can violate the area theorem, regardless of whether the symmetry breaking is explicit or spontaneous Bass:2021acr ; Higgs2014 ; Jeong:2022zea . As indicated in Santos:2021orr ; Sokoliuk:2022llp ; Santos:2023flb ; Santos:2023mee , the interplay between symmetry breaking and residual entropy provides crucial insights into the nature of quantum states in gravitational contexts. Hence, these studies are useful concerning quantum gravity and information theory, particularly within frameworks that account for symmetry breaking.
This work is summarized as follows. In Sec. 2 we present the doubly holographic black string and we extract the conditions that this solution satisfies to become a solution of the Horndeski gravity. In Sec. 3 we calculate the density functional to provide the entanglement thermodynamics in Horndeski gravity, and we extract the entanglement entropy through spatial embedding formalism and Field Theory computation, while in Sec. 4 we present the Page curve behavior. Finally, Sec. 5 provides our conclusions.
2 Horndeski gravity on doubly holographic black string
In order to investigate the structure of the AdS4/BCFT3 correspondence, we start by embedding a three-dimensional Karch-Randall (KR) brane Randall:1999vf ; Karch:2000ct in a four-dimensional black string. The proposal to analyze this scenario within the context of Horndeski gravity allows for an investigation of the effects of modified gravity on the system behavior, particularly regarding thermal states and the degrees of freedom associated with the black hole background. Hence, we consider the following ansatz of a four-dimensional black string, truncated by a Karch-Randall brane, described by the metric
$ds^{2}_{AdS_{4}}=\frac{1}{r^{2}\sin^{2}(u)}\left(-f(r)dt^{2}+dy^{2}+r^{2}du^{2}+\frac{dr^{2}}{f(r)}\right).$ (1)
In our prescription $u\,\epsilon\,[-\infty,\infty]$ and $u\,=-\infty\,\cup\,\infty$ is the asymptotic boundary and the $KR$ brane is embedded at a constant $u=u_{b}$ slice Basu:2023jtf .
In Fig. 3 we present a schematic representation of the black string. The accessible bulk region extends from $u=u_{b}$ to $u=\infty$. As we observe, the geometry in each constant-$u$ slice of the black string is an eternal AdS4 black hole with asymptotic limits. From the overview point of AdS4/BCFT3 correspondence, the dual field theory is a BCFT3 in an AdS3 black hole background with conformal boundary conditions at $r=0$ with the following metric
$ds^{2}_{AdS_{3}}=\frac{1}{r^{2}}\left(-f(r)dt^{2}+dy^{2}+\frac{dr^{2}}{f(r)}\right)\,.$ (2)
Figure 3: Schematic representation of the black string. The $KR$ brane is the point $u=u_{b}$ (red line), the dashed arc at the $r=r_{H}$ represents the black string horizon, and the violet arc is a Hatman-Maldacena (HM) surface Hartman:2013qma . Finally, the island surfaces correspond to the blue curves.
Our scenario incorporates Karch-Randall branes $Q^{L}\,\cup\,Q^{R}$ (i.e. with asymptotically AdS3 geometries) into an asymptotically AdS3 volume described by the following action
$\!\!\!\!\!\!S=\kappa\int_{\mathcal{N}}{d^{4}x\sqrt{-g}\mathcal{L}_{H}}+S^{\mathcal{N}}_{mat}+2\kappa\int_{Q^{L}\,\cup\,Q^{R}}{d^{3}x\sqrt{-h}\mathcal{L}_{bdry}}$ (3)
$\ \ \,+2\int_{Q^{L}\,\cup\,Q^{R}}{d^{3}x\sqrt{-h}\mathcal{L}_{mat}}+2\kappa\int_{Q^{L}\,\cup\,Q^{R},ct}{d^{3}x\sqrt{-h}\mathcal{L}_{ct}}\,,$ (4)
where $\mathcal{N}$ are the two branes form the boundary of a wedge (see Fig. 3) Basu:2023jtf ; Geng:2021mic ; Geng:2023qwm ; Geng:2021hlu ; Geng:2023zhq ; Geng:2022dua ; Saridakis:2007ns ; Kofinas:2014qxa . Furthermore, $S^{\mathcal{N}}_{mat}$ describes ordinary matter that is supposed to be a perfect fluid, and $\mathcal{L}_{mat}$ is a Lagrangian of possible matter fields on $Q^{L}\,\cup\,Q^{R}$. In the above action we have introduced the specific but quite general Horndeski Lagrangians Santos:2021orr ; Santos:2023flb ; Santos:2023mee :
${\cal L}_{H}=(R-2\Lambda)-\frac{1}{2}(\alpha g_{\mu\nu}-\gamma\,G_{\mu\nu})\nabla^{\mu}\phi\nabla^{\nu}\phi\,,$ (5)
$\mathcal{L}_{bdry}=(K-\Sigma)+\frac{\gamma}{4}(\nabla_{\mu}\phi\nabla_{\nu}\phi n^{\mu}n^{\nu}-(\nabla\phi)^{2})K+\frac{\gamma}{4}\nabla_{\mu}\phi\nabla_{\nu}\phi K^{\mu\nu}\,,$ (6)
where $\phi$ is the Horndeski scalar field and $\alpha$ and $\gamma$ are the Horndeski parameters. In the boundary Lagrangian (6), $K_{\mu\nu}=h^{\beta}_{\mu}\nabla_{\beta}n_{\nu}$ is the extrinsic curvature, $h_{\mu\nu}$ is the induced metric and $n^{\mu}$ is the normal vector of the hypersurface $Q^{L}$($Q^{R}$). The traceless contraction of $K_{\mu\nu}$ is $K=h^{\mu\nu}K_{\mu\nu}$, and $\Sigma^{L}$($\Sigma^{R}$) is the boundary tension on $Q^{L}$($Q^{R}$). Finally, ${\cal L}_{ct}$ are boundary counterterms localized on $\mathcal{P}$ Santos:2021orr ; Santos:2023flb ; Santos:2023mee , which is required to be an asymptotic AdS spacetime, and are given by
$\mathcal{L}_{ct}=c_{0}+c_{1}R+c_{2}R^{ij}R_{ij}+c_{3}R^{2}+b_{1}(\partial_{i}\phi\partial^{i}\phi)^{2}+....$ (7)
Assuming that $S^{\mathcal{N}}_{mat}$ is constant, and varying $S^{\mathcal{N}}$ with respect to $g_{\alpha\beta}$ and $\phi$, we obtain the field equations
${\cal E}_{\mu\nu}[g_{\mu\nu},\phi]$ $=$ $G_{\mu\nu}+\Lambda g_{\mu\nu}-\frac{\alpha}{2}\left(\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\nabla_{\lambda}\phi\nabla^{\lambda}\phi\right)$ (8)
-- $\frac{\gamma}{2}\left(\frac{1}{2}\nabla_{\mu}\phi\nabla_{\nu}\phi R-2\nabla_{\lambda}\phi\nabla_{(\mu}\phi R^{\lambda}_{\nu)}-\nabla^{\lambda}\phi\nabla^{\rho}\phi R_{\mu\lambda\nu\rho}\right)$
-- $\frac{\gamma}{2}\left(-(\nabla_{\mu}\nabla^{\lambda}\phi)(\nabla_{\nu}\nabla_{\lambda}\phi)+(\nabla_{\mu}\nabla_{\nu}\phi)\Box\phi+\frac{1}{2}G_{\mu\nu}(\nabla\phi)^{2}\right)$
$+$ $\frac{\gamma g_{\mu\nu}}{2}\left(-\frac{1}{2}(\nabla^{\lambda}\nabla^{\rho}\phi)(\nabla_{\lambda}\nabla_{\rho}\phi)+\frac{1}{2}(\Box\phi)^{2}-(\nabla_{\lambda}\phi\nabla_{\rho}\phi)R^{\lambda\rho}\right),$
${\cal E}_{\phi}[g_{\mu\nu},\phi]$ $=$ $\nabla_{\mu}\left[\left(\alpha g^{\mu\nu}-\gamma G^{\mu\nu}\right)\nabla_{\nu}\phi\right]\,.$ (9)
The no-hair theorem requires that the square of the radial component of the conserved current vanishes identically without restricting the radial dependence of the scalar field, which implies
$\alpha g_{rr}-\gamma G_{rr}=0\,,$ (10)
and from this condition we find ${\cal E}_{\phi}[g_{rr},\phi]=0$ Santos:2021orr ; Santos:2023flb ; Santos:2023mee . Without loss of generality we consider $\phi=\phi(r)$ and we define $\phi^{{}^{\prime}}(r)\equiv\psi(r)$. As it can be shown, the equations ${\cal E}_{\phi}[g_{rr},\phi]={\cal E}_{rr}[g_{rr},\phi]=0$ are satisfied, and therefore we can calculate the horizon functions $f(r)$ and $\psi(r)$ as (for the AdS4 radius we set 1)
$f(r)$ $=$ $1+Cr^{3}+\left(-3+\frac{\alpha}{\gamma}\right)\csc^{2}(u),$ (11)
$\psi^{2}(r)$ $=$ $-\frac{4(\alpha+\gamma\Lambda)}{\alpha\gamma r^{2}A(u)\sin^{2}(u)}\frac{1}{f(r)}.$ (12)
$A(u)$ $=$ $2\alpha-5\gamma-\gamma\cos(2u),$ (13)
where equations (11)-(12) are found though (10) with $r>0$, $0<u<\pi$ and $x,y\in\mathbb{R}$. The blackening factor $f(r)$ gives a black hole on each constant-$u$ slice, and one such slice will be the KR brane Randall:1999vf ; Karch:2000ct . Finally, replacing (11) and (12) in Horndeski equations of motion, we find the following conditions
$\frac{(\beta_{0}+1)}{A(u)}=0$ (14)
$\frac{(\beta_{0}+1)}{(\alpha-\gamma)}-\frac{(\beta_{0}+1)}{A(u)}=0,$ (15)
[…]
$\left[K_{\alpha\beta}-h_{\alpha\beta}(K-\Sigma)+\frac{\gamma}{4}H_{\alpha\beta}\right]_{Q^{L}\,\cup\,Q^{R}}=0$ (16)
$H_{\alpha\beta}\equiv(\nabla_{\mu}\phi\nabla_{\nu}\phi\,n^{\mu}n^{\nu}-(\nabla\phi)^{2})(K_{\alpha\beta}-h_{\alpha\beta}K)-(\nabla_{\mu}\phi\nabla^{\mu}\phi)h_{\alpha\beta}K\,,$ (17)
where the defect is given
$y{{}^{\prime}}(r)$ $=$ $\frac{(\Sigma)}{\sqrt{1-\dfrac{\xi}{1-\left(\frac{r}{r_{h}}\right)^{2}}-(\Sigma)^{2}\left(1-\left(\frac{r}{r_{h}}\right)^{2}\right)}}\,;\,\xi=-\frac{1}{2}\frac{\alpha+\gamma\Lambda}{\alpha}.$ (18)
We mention here that for Horndeski gravity we have a restriction on the parameters of the theory due to the lack of a consistent solution within this theory. On the other hand, for the case of Dvali-Gabadadze-Porrati (DGP) scenario Dvali:2000hr the entanglement island is not affected by the deformation of its configuration, thus providing constraints on the DGP model through holographic consistency. Heisenberg:2018vsk ; Geng:2023qwm .
The analysis of the island region surrounding the black hole horizon within the framework of Horndeski gravity can be useful for the study of quantum information dynamics and entanglement Geng:2021mic . The observation that the defect cannot be accessed from the island without traversing its complement reveals the topological and geometrical complexities inherent in these scenarios, as illustrated in Fig. 3. Finally, note that the size of the island is contingent upon the choice of brane angle $u_{b}$.
Figure 4: The relevant Ryu-Takayanagi (RT) surfaces with $\alpha=-8/3$ and $\Lambda=-3$ for different values of the $\gamma$-Horndeski parameter, namely $\gamma=-0.1$ (solid), $\gamma=-0.2$ (dashed), $\gamma=-0.3$ (dot dashed), and $\gamma=-0.4$ (thick). The islands are always attached to the brane inside the atoll; they start at the point $r_{L}$ and extend to the horizon of the black hole. Decreasing $\gamma$ reduces the size of the atoll, which shrinks towards the horizon, while increasing the size of $R$, limited by $r_{L}$, pushes the anchors towards the black hole horizon. Finally, increasing the angle $u_{b}$ of the brane while keeping the Horndeski parameters fixed can push the anchors toward the defect.
Let us now examine the different classes of Ryu-Takayanagi (RT) surfaces Geng:2021mic in Horndeski gravity. The formation of the largest possible island, when the island surface originates at an anchor point $r_{L}$ and extends to the defect (see Fig. 4), reveals the role of boundary conditions and spatial geometry in affecting the entanglement structure. When $r_{L}$ coincides with the defect, resulting in the atoll enveloping the brane entirely, it marks a significant scenario where entanglement properties are maximized. This suggests a profound connection between the island geometry, the defect location, and the resultant entanglement entropy.
At the boundary $r=0$ with $\Sigma=\cos(u_{b})$, Eq. (18) becomes
$y(r)=u_{0}+\frac{r\cos(u_{b})}{\sqrt{-\xi}},$ (19)
and considering for instance $u_{b}=\pi/2$ or $23\pi/$ 2, we acquire $y(r)=u_{0}={\rm constant}$. The intersection of the branes at the boundary $r=0$ marks their critical role in defining the asymptotic structure of spacetime. The presence of positive tension on both branes, denoted as $\Sigma^{L}$($\Sigma^{R}$), indicates their influence on the geometry, effectively constraining the central bulk region between them and determining the overall gravitational configuration. As the branes approach their corresponding limits, the left brane at $y=0$ and the right brane at $y=\pi$ establish natural boundary conditions for the dual field theory. The positive tension suggests that the branes exert a repulsive force against the bulk, potentially amplifying curvature effects in their vicinity and impacting the properties of possible quantum fields residing on the branes.
The exploration of the implications of Dirichlet and Neumann boundary conditions is important for understanding the construction of RT surfaces, and can lead to different physical interpretations regarding the flow of information and the stability of entangled states in the boundary theory. Nevertheless, the different angles of the brane, particularly below the critical angle (see Fig. 4), introduces more complications. In particular, the critical angle can mark a transition in the behavior of the RT surfaces, possibly affecting their stability and the nature of the entanglement islands, and surfaces that connect to the brane at angles below this critical threshold might exhibit different entanglement properties than those that do not.
Figure 5: $Q$ boundary profile for the AdS3 black hole within Horndeski gravity considering the values $\theta^{\prime}=2\pi/3$, $u_{b}=\pi-\theta^{\prime}$, $\alpha=-8/3$, $\Lambda=-3$, and with $\gamma=-0.1$ (solid), $\gamma=-0.2$ (dashed), $\gamma=-0.3$ (dot dashed), and $\gamma=-0.4$ (thick). The dashed parallel vertical lines represent the ultra-violet (UV) solution (18), while the region between curve $Q$ negative and positive branches represents the bulk $\mathcal{N}$ Santos:2021orr .
In Fig. 5 we depict the $Q$ boundary profile for the AdS3 black hole within Horndeski gravity. As we observe, the internal dynamics of the KR-braneworld and the presence of Horndeski gravity complicates the behavior of the entanglement islands Brito:2018pwe ; Santos:2023eqp . In particular, the formation of the largest possible island, starting from a critical anchor and ending at the fault, serves as a key feature in determining the entanglement structure of the system. The critical anchor represents a point from which the entanglement island can expand, while the fault acts as a boundary restricting this expansion. This relation emphasizes the importance of geometric considerations in entanglement dynamics, showing how the shapes and connections of these regions directly affect the overall entanglement entropy Geng:2021mic . Hence, atolls are crucial to understand the phase structure of the system, since their formation and characteristics can lead to different phases of entanglement, which may be associated with varying behaviors of the entanglement entropy.
3 Entanglement entropy and Hatman-Maldacena surface
In this section we present the entanglement entropy to the subregion $\mathcal{R}$, $\bar{\mathcal{R}}$ and defect, which leads to the island on the brane. In general, it is possible to find extremal surfaces $\Omega$ that satisfy the holonomy constraint like $\partial\,\Omega=\partial\mathcal{\mathcal{R}}\cup\partial\bar{\mathcal{R}}$. In scenarios involving defects or branes, the presence of an island can complicate the calculation of entanglement entropy. The interaction between the two islands is essential for ensuring that the entanglement entropy adheres to unitarity. As the system evolves, the total entropy reflects information conservation. The dominance of an island surface over an Hatman-Maldacena (HM) Hartman:2013qma surface at an initial time indicates that the entanglement dynamics can shift, revealing the non-trivial evolution of entanglement in these scenarios. Page time is particularly relevant here, as it marks the moment when the entanglement entropy transit from being dominated by the HM surface to the island surface. The dependence of Page time on parameters such as $u_{b}$, $\alpha$, $\gamma$, and the sizes $r_{L}$ and $r_{R}$, indicates the richness of the model. Each parameter plays a role in determining the degrees of freedom in the defect and the size of the regions of interest Takayanagi:2011zk . Finally, the observation that Page time is approximately proportional to the difference in area between the RT surfaces at $t=0$, reveals a geometric aspect of the entanglement dynamics, and this area difference reflects the competing contributions to the entanglement entropy and significantly determines how the entropy evolves.
Before presenting the entanglement entropy and Hatman-Maldacena surface in Horndeski’s scenario. We will present entanglement entropy calculation in the Karch-Randall braneworld in Sec. 3.1, it emerges naturally as presented by Geng:2024xpj . Based on this scenario, we can extend the calculation of entanglement entropy to the Horndeski scenario Sec. 3.2.
3.1 Entanglement entropy calculation in the Karch-Randall braneworld
The Karch-Randall braneworld scenario provided a natural setting to study Hawking radiation from a black hole using holographic tools Randall:1999vf ; Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp . However, quantum effects were an open problem until recently, leaving a gap that needs to be filled. This gap was filled in the work Geng:2024xpj where the ambient space is (2+1)-dimensional, for which explicit calculations can be made in each configuration description.
As discussed and shown by Geng:2024xpj , the Karch-Randall braneworld model involves embedding a subcritical brane within an ambient asymptotically AdSd+1 spacetime, denoted as $\mathcal{N}_{d+1}$. In this framework, the ambient spacetime is governed by classical Einstein gravity, while the brane is a codimension-one hypersurface, $\mathcal{N}_{d}$. The action governing this system is given by
$S=-\frac{1}{16\pi G_{d+1}}\int_{\mathcal{N}_{d+1}}d^{d+1}x\sqrt{-g}(R-2\Lambda)-\frac{1}{8\pi G_{d+1}}\int_{\mathcal{N}_{d}}d^{d}x\sqrt{-h}(K-T)\,,$ (20)
The boundary conditions for the bulk metric in the Karch-Randall braneworld model are of the Neumann type near the brane and the Dirichlet type near the remaining asymptotic boundary Geng:2023qwm ; Geng:2021mic . This implies nonzero metric fluctuations near the brane, while fluctuations vanish near the asymptotic boundary. In this framework, the brane represents a gravitational AdSd+1 spacetime, whereas the asymptotic boundary acts as a $d$-dimensional nongravitational bath Geng:2021iyq ; Geng:2020fxl .
Two sets of equations of motion govern the system:
  • Einstein’s equations, which ensure the vanishing of the bulk variation of the action (20), and the brane embedding equations, vanishing to the brane for the variation of the same action;
  • Einstein’s equations determine the background geometry of the ambient spacetime. In contrast, the brane embedding equations dictate how the brane is embedded within this spacetime, thereby defining the induced geometry on the brane Santos:2021orr ; Geng:2024xpj .
For the bulk description is asymptotically AdS3 with Karch-Randall braneworld Karch:2000ct , this description is associated to BCFT2; we want to calculate the subregion entanglement entropy SR of the thermal field dual state Geng:2021mic . To study the thermal field double states $|TFD\rangle$, one intrudes two Karch-Randall branes (see Fig. 6) into the spacetime of the ambient BTZ black hole; the BCFT2 are in this double state. In this boundary description, the two BCFT2 live in strips with boundary conditions corresponding to Cardy states Cardy:2004hm .
Figure 6: Configuration with two branes in the BTZ background where the gray regions represent cut off by the branes.
In the setup illustrated in Fig. 7 Geng:2024xpj , we consider a model involving two eternal black holes in (1+1)-dimensions, which are coupled through a thermal bath (see Fig. 7-a). This configuration is analyzed within the framework of Boundary Conformal Field Theory (BCFT), where the boundaries are characterized by specific boundary conditions associated with Cardy states (refer to Fig. 7-b) Ryu:2006bv ; Cardy:2004hm ; Geng:2024xpj . This setup provides a novel perspective on the interaction dynamics between black holes and their environments, offering insights into the underlying quantum gravitational processes.
Figure 7: The figure shows in (a) Penrose diagram of the intermediate description for two branes at the bottom of the BTZ black hole; we have two ($1+1$)-dimensional black holes, which are coupled to each other through strip-shaped baths (the green shaded region). We identify the two outer edges in red. It can be seen that the black hole singularities are orbifold singularities inherited from the BTZ black holes. In part of frame (b), we have a description of the boundary of the configuration with two branes at the bottom of the BTZ black hole.
Through the holographic dictionary of the Karch-Randall braneworld, one can perform a mapping of the replicated boundary partition function into a Euclidean gravitational path integral as follows:
$Z_{\text{BCFT}}[\mathcal{N}^{(n)}]=\int_{\partial\mathcal{N}^{(n)}_{d+1}=\mathcal{N}^{(n)}}D[g]e^{-S}=Z_{\text{grav}}[\mathcal{N}^{(n)}]\,,$ (21)
[…]
$=$ $\lim_{n\rightarrow 1}\frac{1}{1-n}\frac{Z_{\text{grav}}[\mathcal{N}^{(n)}]}{Z_{\text{grav}}[\mathcal{N}^{(0)}]^{n}}\,,$ (23)
[…]
Figure 8: The diagram illustrates the Euclidean path integral formulation with an initially prepared ground state, denoted as $|0\rangle$. Within the dashed interval, we identify the subregion $\mathcal{R}$. The reduced density matrix operator for this subregion, $\hat{\rho}_{\mathcal{R}}$, is derived from the full density matrix $\hat{\rho}=|0\rangle\langle 0|$ by integrating over all possible field configurations outside of $\mathcal{R}$. This approach allows us to isolate the contributions specific to the subregion, providing insights into the entanglement properties and quantum correlations present in the system.
Through the half Euclidean plane with the Cardy boundary-$\mathcal{N}$ Ryu:2006bv ; Cardy:2004hm ; Geng:2024xpj , the trace of the $n-th$ power of the normalized reduced density matrix operator is given by
$tr\hat{\rho}_{\bar{R}}^{n}=\frac{Z_{BCFT}[\mathcal{N}^{(n)}]}{Z^{n}_{BCFT}[\mathcal{N}]}\,.$ (25)
This expression is the key point to prove the Ryu-Takayanagi Ryu:2006bv conjecture in the AdS/BCFT correspondence scenario using gravitational path integral. Thus, with this, we can conclude that
$S_{R}$ $=$ $\lim_{n\rightarrow 1}\frac{1}{1-n}\frac{e^{-nS[g^{BTZ}]+\frac{A(\gamma)}{4G_{d+1}}\frac{1-n}{n}]}}{e^{-nS[g^{BTZ}]}}\,,$ (26)
$=$ $\frac{A(\gamma)}{4G_{d+1}}\,.$ (27)
where $g^{EBTZ}$ is the Euclidean BTZ metric. The phases of the Karch-Randall brane within the manifold $\mathcal{N}_{d+1}^{(n)}$ Geng:2021mic ; Geng:2022dua .
With this prescription, we can identify two distinct phases:
  • First phase: characterized by $n$-disconnected components and another by a single component that connects the Cardy boundaries of the boundary replica manifold $\mathcal{N}_{d+1}^{(n)}$ Cardy:2004hm . In this phase, the conical singularity $\gamma$ remains in the bulk, linking the two $\partial\mathcal{R}$ boundaries;
  • Second phase: $\gamma$ extends to the brane, with $\mathbb{Z}_{n}$ fixed points present, resulting in two disconnected components Tonni:2010pv that connect the boundaries of $\partial\mathcal{R}$ to the brane Geng:2023qwm ; Geng:2021hlu .
This dual-phase behavior provides new insights into the geometric and topological properties of the Karch-Randall brane, offering potential implications for understanding brane dynamics in higher-dimensional spaces Geng:2022dua . However, we have that
$S(R)=\min(\frac{A_{c}}{4G_{d+1}},\frac{2A_{dc}}{4G_{d+1}})\,,$ (28)
This is the Ryu-Takayanagi conjecture Ryu:2006bv . Note that $\gamma_{dc}$ must also have its area minimized over its possible endpoints on the brane; this corroborates our saddle point approximation. This discussion shows that the Ryu-Takayanagi conjecture is proven; one can see that, indeed, the emergence of the disconnected Ryu-Takayanagi surface is due to the connected phase of the Karch-Randall brane in the replica path integral Basu:2023jtf .
3.2 Derivation of the functional entanglement entropy for Horndeski gravity
We must derive the functional area for Horndeski gravity’s entanglement entropy to study the holographic entanglement entropy. Considering the Horndeski Lagrangian ${\cal L}_{H}$, and following the technique detailed in previous works of DosSantos:2022exb ; Caceres:2017lbr ; Feng:2015oea ; Wald:1993nt ; Iyer:1994ys , which consider:
  • The Lagrangian-${\cal L}_{H}$ of Horndeski theory does not include quadratic or higher-order curvature terms;
  • Incorporates a Wald-type term Feng:2015oea , which is important in developing holographic entanglement entropy in arbitrary gravity theories.
This approach facilitates the application of the replica trick, a method essential for deriving holographic entanglement entropy. Through the microscopic definition of entanglement entropy Ryu:2006ef ; Caceres:2017lbr :
$S_{EE}=-Tr(\rho\log(\rho))$ (29)
Here, $\rho$ is the reduced density matrix of a subregion of a time slice of the boundary. Here, the idea is to find the entanglement entropy in the subregion $\mathcal{R}$, $\bar{\mathcal{R}}$. For this, we apply the replica trick through the $n^{th}$ Rényi entropy, defined by:
$S_{n}=-\frac{1}{n-1}\log{\mathrm{Tr}{(\rho^{n})}}.$ (30)
Here the entanglement entropy is the analytical continuation of $S_{n}$ as $n\rightarrow 1$: $S_{EE}=\lim_{n\rightarrow 1}S_{n}$. From the overview point of path integral, $\rho$, can be express $S_{n}$ in terms of the partition function $Z_{n}$ considering an appropriate n-sheeted Riemann surface $M_{n}$ as:
$S_{n}=-\frac{1}{n-1}\left(\log{Z_{n}}-n\log{Z_{1}}\right),$ (31)
in the above equation $Z_{1}$ is the original partition function. Besides, using the basic holographic relation $Z_{CFT}=e^{-S_{bulk}}$ between the field theory partition function $Z_{CFT}$ and the bulk action for an appropriate bulk geometry. With this and through the prescription of Caceres:2017lbr ; Feng:2015oea , we can derive the area functional for Entanglement Entropy for Horndeski gravity ($S_{EEH}$):
$S_{EEH}=\int_{\mathcal{N}}{d^{d}x\sqrt{\eta}\frac{\partial{\cal L}_{H}}{\partial R_{\mu\nu\rho\sigma}}\epsilon_{\mu\nu}\epsilon_{\rho\sigma}},$ (32)
where $\epsilon_{\mu\nu}$ and $\epsilon_{\rho\sigma}$ denotes the bi-normal to $\mathcal{N}$ Feng:2015oea ; Wald:1993nt ; Iyer:1994ys . We can derive Horndeski Lagrangian with respect to the Riemann tensor as:
$\frac{\partial{\cal L}_{H}}{\partial R_{\mu\nu\rho\sigma}}$ $=$ $-\frac{\kappa}{32\pi}(g^{\mu\rho}g^{\nu\sigma}-g^{\nu\rho}g^{\mu\sigma})-\frac{\gamma}{128\pi}[g^{\mu\rho}\phi^{,\nu}\phi^{,\sigma}-g^{\nu\rho}\phi^{,\mu}\phi^{,\sigma}+g^{\nu\sigma}\phi^{,\mu}\phi^{,\rho}-g^{\mu\sigma}\phi^{,\nu}\phi^{,\rho}$ (33)
-- $(g^{\mu\rho}g^{\nu\sigma}-g^{\nu\rho}g^{\mu\sigma})\phi^{,\lambda}\phi_{,\lambda}]$
Here in our work $\phi^{,\alpha}=g^{\alpha\beta}\phi_{,\beta}$. For the metric in complex coordinates like:
$ds^{2}=dzd\bar{z}+\eta_{ij}dy^{i}dy^{j}$ (34)
Upon substituting the metric components into (33), we have
$\frac{\partial{\cal L}_{H}}{\partial R_{z\bar{z}z\bar{z}}}=\frac{\kappa}{8\pi}-\frac{\gamma}{32\pi}\left(\phi^{,\lambda}\phi_{,\lambda}-\phi^{,z}\phi^{\bar{,z}}\right)$ (35)
Upon further expanding $\phi^{,\lambda}\phi_{,\lambda}=\phi^{,z}\phi^{,\bar{z}}+\eta_{ij}\phi^{,i}\phi^{,j}$, this further simplifies to:
$\frac{\partial{\cal L}_{H}}{\partial R_{z\bar{z}z\bar{z}}}=\frac{\kappa}{8\pi}-\frac{\gamma}{32\pi}\eta_{ij}\phi^{,i}\phi^{,j}$ (36)
Thus, we have that the functional for holographic entanglement entropy then read Caceres:2017lbr ; Feng:2015oea :
$S_{EEH}=\int{d^{d}x\sqrt{\eta}\frac{\partial{\cal L}_{H}}{\partial R_{z\bar{z}z\bar{z}}}}=\frac{\kappa}{4}\int{d^{d}x\sqrt{\eta}\left[1-\frac{\gamma}{4\kappa}\eta_{ij}\phi^{,i}\phi^{,j}\right]}$ (37)
We can note that comparing our result with the Ryu-Takayanagi formula Ryu:2006ef , the correction for the area is due to the term $\gamma\,G_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi$ where for $\gamma=0$ we recover the usual Einstein case. Considering the metric (1) and (37), we can derive the area functional $\mathcal{A}(\partial\mathcal{\mathcal{R}}\cup\partial\bar{\mathcal{R}})$ to the subregion $\mathcal{R}$ and $\bar{\mathcal{R}}$ for Horndeski gravity, which is given by
$S_{EEH}=\frac{\mathcal{A}}{4G}$ (38)
[…]
$\chi=1+\frac{3(\alpha+\gamma\Lambda)}{\alpha\,A(u)}.$ (40)
Now, the Euler-Lagrange equation for the action (39) reads
$r{{}^{\prime\prime}}=-rf(r)+2\cot(u)r{{}^{\prime}}+\frac{2\cot(u)r{{}^{\prime}3}}{r^{2}f(r)}-\frac{r{{}^{\prime}}\chi{{}^{\prime}}}{\chi}-\frac{r{{}^{\prime}3}\chi{{}^{\prime}}}{r^{2}f(r)\chi}+\frac{r{{}^{\prime}2}}{r}.$ (41)
[…]
$\chi=\frac{2-\beta_{0}(1-\beta_{0})}{2(1-\beta_{0})}=constant.$ (42)
Hence, we recover the result of Geng:2021mic , namely
$r{{}^{\prime\prime}}=-rf(r)+2\cot(u)r{{}^{\prime}}+\frac{2\cot(u)r{{}^{\prime}3}}{r^{2}f(r)}+\frac{r{{}^{\prime}2}}{r}.$ (43)
Figure 9: The critical angle and ancor with $\alpha=-8/3$ and $\Lambda=-3$ for different values of the Horndeski parameter $\gamma$, namely $\gamma=-0.1$ (solid), $\gamma=-0.2$ (dashed), $\gamma=-0.3$ (dot dashed), and $\gamma=-0.4$ (thick). The critical angle is a monotonically increasing function of the number of spatial dimensions $d=4$, thus the behavior of the critical anchor that defines the beginning of the atoll for the black string also assumes different values (see Fig. 4). Finally, the critical anchor increases monotonically with the brane angle and decreases as it coincides with the defect at the critical angle.
As we can see, according to Geng:2021mic the boundary terms in the variation of $\mathcal{A}$ vanish by imposing boundary conditions on $\Sigma$. This occurs since a Dirichlet condition is imposed on the conformal boundary and a Neumann condition Geng:2020fxl . However, the angles are deformed by the Horndeski parameters and must anchor on the brane at right angles; they will be right angles only for $\alpha=0$ and $\gamma=0$. Thus, the Euler-Lagrange equation with Horndeski parameters $\alpha$ and $\gamma$ leads to nontrivial physics below the critical angle we explore. In Fig. 9 we analyze how the critical angle depends on $\alpha$ and $\gamma$, giving the trajectories of RT surfaces becoming circular geodesics that exist for all brane angles.
3.3 Hartman-Maldacena entanglement entropy
We use the spatial embedding formalism to compute the Hartman-Maldacena (HM) entanglement entropy Hartman:2013qma . Thus, the geometry is embedded as a codimension-one sub-manifold of a four-dimensional Minkowski space:
$ds^{2}=\eta_{ab}dX^{a}dX^{b};\,\eta_{ab}=diag(1,1,-1,-1),$ (44)
where the embedding equation is $X_{a}X^{a}=1$. For convenience, we impose the re-parameterization $sin^{-1}(u)=\cosh(\rho)$ for the metric (1), and considering the fact that a dual field theory is a BCFT2 on an AdS2 black hole background with conformal boundary conditions at $r=0$, we obtain
$ds^{2}_{AdS_{4}}=\frac{\cosh(\rho)}{r^{2}}\left(-f(r)dt^{2}+\frac{dr^{2}}{f(r)}\right)+d\rho^{2}.$ (45)
To simplify our results, we rewrite $f(r)$ as
$f(r)=\omega\left(1-\frac{r^{2}}{r^{2}_{h}}\right);\,\omega=1+\left(\frac{\alpha-3\gamma}{\gamma}\right)\cosh^{2}(\rho),$ (46)
and therefore the metric (45) can be recovered using the following parameterization of the embedding equation
$X_{0}=\frac{2r_{h}-r}{r}\cosh(\rho)$ (47)
$X_{1}=\frac{2r_{h}}{r}\sqrt{\omega\left(1-\frac{r^{2}}{r^{2}_{h}}\right)}\sin\left(\frac{2\pi\,t}{\beta}\right)\cosh(\rho)$ (48)
$X_{2}=\frac{2r_{h}}{r}\sqrt{\omega\left(1-\frac{r^{2}}{r^{2}_{h}}\right)}\cos\left(\frac{2\pi\,t}{\beta}\right)\cosh(\rho)$ (49)
$X_{3}=\sinh(\rho),$ (50)
with $\beta=4\pi\,r_{h}=2\omega/T$ the inverse Hawking temperature.
The idea behind the spatial embedding formalism is that we can calculate the Hartman-Maldacena surface area without solving the minimum area (geodesic) differential equation. Using this embedding, the length $l$ can be computed with the following coordinates ($X_{0},\,X_{1},\,X_{2},\,X_{3}$) and ($X^{\prime}_{0},\,X^{\prime}_{1},\,X^{\prime}_{2},\,X^{\prime}_{3}$):
$l=\cosh^{-1}(X_{0}X^{\prime}_{0}X_{1}X^{\prime}_{1}-X_{2}X^{\prime}_{2}-X_{3}X^{\prime}_{3}).$ (51)
As shown in the schematic diagram of Fig. 2, while the island surface begins at the bipartition and ends at the KR brane, we can observe that the HM surface Hartman:2013qma passes through an Einstein bridge-Rosen and ends at the right bipartition in the double thermal field on the right side. In our diagram, the left and right bipartitions are located at $(u;\rho)=(u_{L};1)$ and $(u;\rho)=(u_{R};1)$. As discussed in Geng:2022dua , one can introduce a regularization parameter $\rho_{\epsilon}$, which can be set to 1; our bipartitions are on the asymptotic boundary. We can note that in order to obtain a correct bipartition we must take the time coordinate $t\to-t+\frac{i\beta}{2}$. This corresponds to the reversal of the Killing vector field similarly to the time on the other side of the black string horizon Geng:2022dua (see Fig. 2).
We proceed by introducing the quantities
$\Delta_{L}=r_{h}-r_{L},\,\Delta_{R}=r_{h}-r_{R}.$ (52)
Hence, the HM surface can be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\mathcal{A}_{HM}=\chi\cosh^{-1}(-X^{L}_{0}X^{R}_{0}-X^{L}_{1}X^{R}_{1}+X^{L}_{2}X^{R}_{2}+X^{L}_{3}X^{R}_{3})$
$=\chi\cosh^{-1}\left[\frac{(2r_{h}-r_{L})(2r_{h}-r_{R})+4r_{h}\omega(\rho_{\epsilon})\sqrt{\Delta_{L}\,\Delta_{R}}\,\cosh\left(\frac{2\pi\,t}{\beta}\right)}{r_{L}r_{R}}\cosh(\rho_{\epsilon})-\sinh^{2}(\rho_{\epsilon})\right]\!\!.$ (53)
Using the hyperbolic trigonometric identities, we acquire
$S_{HM}$ $=$ $\frac{\mathcal{A}_{HM}}{4G}$ (54)
$=$ $\frac{c}{6}log\,\left[\frac{r_{h}}{r_{L}r_{R}}\left(\Delta_{L}+\Delta_{R}+2\omega(\rho_{\epsilon})\sqrt{\Delta_{L}\,\Delta_{R}}\,\cosh\left(\frac{4\pi\,t}{\beta}\right)\right)\right]+\frac{c}{3}\rho_{\epsilon},$ (55)
where $c=3\chi/2G$ is the central charge-like for Horndeski gravity DosSantos:2022exb . Thus, considering a particular case with $r=r_{L}=r_{R}$, and using $\Delta_{r}=r_{h}-r$, we have
$S_{HM}=\frac{c}{6}log\,\left[\frac{2r_{h}\Delta_{r}}{r^{2}}\left(1+\omega(\rho_{\epsilon})\sqrt{\Delta_{L}\,\Delta_{R}}\,\cosh\left(\frac{4\pi\,t}{\beta}\right)\right)\right]+\frac{c}{3}\rho_{\epsilon}.$ (56)
We can now obtain the result with the pair of minimal island surfaces that cross from the bipartitions $r_{L}$ and $r_{R}$ to the corresponding physical branes location at $\rho=\rho_{*}$. In particular, we can schematically represent $\mathcal{A}_{island}$ as:
[…]
In this case, the entanglement entropy calculated by the island surface is given by
$S_{island}=\frac{\mathcal{A}_{island}}{4G}=-\frac{c}{3}\rho_{*}+\frac{c}{3}\rho_{\epsilon},$ (59)
where the prescriptions of RT Ryu:2006bv show that the entanglement entropy is the minimum of $S_{HM}$ and $S_{island}$, i.e. $S=min(S_{HM},S_{island})$.
3.4 Field theory calculation
In this subsection we calculate the entanglement entropy between subsystem $A$ and its complement. This calculation is equivalent to calculating a two-point function $\langle\Phi_{n}(r_{R},t)\,\Phi_{n}(r_{L},t)\rangle$ of the twist operator fields $\Phi_{n}(r,t)$ inserted in the two bipartition points $r_{L}$ and $r_{R}$. Hence, since the field theory lives on a curved background, in order to perform the entanglement entropy calculation on the field theory side we need the appropriate geometry Geng:2022dua ; Doroudiani:2019llj .
We can calculate the entanglement entropy of subsystem $A$ at the boundary ($S_{bdry}$) and in the bulk ($S_{bulk}$) channels, using the conformal transformations Basu:2023jtf ; Geng:2022dua :
$ds^{2}=\frac{1}{r^{2}}\left[-\omega\left(1-\frac{r}{r_{h}}\right)dt^{2}+\frac{dr^{2}}{\omega\left(1-\frac{r}{r_{h}}\right)}\right]$ (60)
$ds_{conformal}^{2}=\Omega^{2}(r_{*})\left(-dt^{2}+dr^{2}\right)$ (61)
$\Omega^{2}(r_{*})=\frac{1}{r}\sqrt{\omega\left(1-\frac{r}{r_{h}}\right)}$ (62)
$r_{*}=-r_{h}log\left[\omega\left(1-\frac{r}{r_{h}}\right)\right],$ (63)
with Conformal boundary (see Fig. 10)
$z=r_{*}+i\tau\ ;\ \overline{z}=r_{*}-i\tau$ (64)
$ds^{2}=\Omega^{2}(r_{*})dzd\overline{z}=\Omega^{2}(r_{*})e^{-\frac{r_{*}}{r_{H}}}d\sigma d\overline{\sigma}.$ (65)
The twist fields $\Phi_{n}$ inserted at the two bipartition point $r_{L}$ and $r_{R}$ satisfy
[…]
Finally, we find
$S_{bdry}=2\,log(g_{b})+\frac{c}{3}\log\left(\frac{2}{\epsilon}\right)$ (67)
$S_{bulk}=\frac{c}{6}log\,\left[\frac{r_{h}}{r_{L}r_{R}}\left(\Delta_{L}+\Delta_{R}+2\omega(\rho_{\epsilon})\sqrt{\Delta_{L}\,\Delta_{R}}\,\cosh\left(\frac{4\pi\,t}{\beta}\right)\right)\right]+\frac{c}{3}\log\left(\frac{2}{\epsilon}\right),$ (68)
with $\Delta_{L}=r_{h}-r_{L}$, $\Delta_{R}=r_{h}-r_{R}$ and where $\epsilon$ is the UV cutoff.
Figure 10: Schematic map of a field theory on the curved geometry of a flat background, using the conformal transformations.
According to Santos:2021orr , we can calculate the Horndeski gravity boundary entropy of the boundaries $Q_{L}$ and $Q_{R}$ for the two bipartitions as
$S^{L}_{bdry}=\frac{2\,\Delta y_{Q_{L}}}{3r_{L}}\left(1-\frac{\xi}{8}\right)-\frac{\xi\,b(u_{b})}{3r^{2}_{L}}\left(1-\frac{\xi}{4}\right)-\frac{\xi\,h(u_{b})\cot(u_{b})}{r^{2}_{L}}+\frac{\xi\,q(u_{b})}{3}$ (69)
$S^{R}_{bdry}=\frac{2\,\Delta y_{Q_{R}}}{3r_{R}}\left(1-\frac{\xi}{8}\right)-\frac{\xi\,b(u_{b})}{3r^{2}_{R}}\left(1-\frac{\xi}{4}\right)-\frac{\xi\,h(u_{b})\cot(u_{b})}{r^{2}_{R}}+\frac{\xi\,q(u_{b})}{3},$ (70)
with
$b(u_{b})$ $=$ $\cos(u_{b})\tan^{-1}\left(\frac{1}{\sin(u_{b})}\right)+\cot(u_{b})\left(\frac{1+\cos^{2}(u_{b})\cot^{2}(u_{b})}{\sin^{2}(u_{b})}\right)\,$ (71)
$h(u_{b})$ $=$ $-\frac{(1+\pi/2)}{2\sin(u_{b})}+\frac{\cot^{3}(u_{b})\cos^{2}(u_{b})}{(1+\cos^{2}(u_{b}))}\tanh^{-1}\left(\frac{\sqrt{2}\cos(u_{b})}{\sqrt{1+\cos^{2}(u_{b})}}\right)$
-- $\frac{(1+\cos^{2}(u_{b})+3\cos^{4}(u_{b})-3\cos^{6}(u_{b}))}{3\sin^{5}(u_{b})(1+\cos^{2}(u_{b}))}\,,$
$q(u_{b})$ $=$ $\left(\frac{1}{4}-\cos^{3}(u_{b})\right)\cot(u_{b})\csc(u_{b})\,.$
Hence, if the subsystems are considered large and far from the boundary then we obtain
$S^{L}_{bdry}=\frac{\xi\,q(u_{b})}{3}$ (72)
$S^{R}_{bdry}=\frac{\xi\,q(u_{b})}{3}.$ (73)
This result expresses a significant fact about the information process in black holes (specifically the geometry of the AdS4 black string with a Karch-Randall brane Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp ). Note that this residual information implies that even if the black hole evaporates entirely from the point of view of classical entropy, we still have information being emitted Susskind:2014rva ; Brown:2015bva ; Brown:2015lvg ; Susskind:2018fmx ; Brown:2018bms ; Brown:2017jil ; Brown:2019whu ; Brown:2022rwi ; Santos:2021orr ; Santos:2023mee ; Santos:2024zoh . In this case $S^{L}_{bdry}$ and $S^{R}_{bdry}$ are holographically entangled and are estimated as the minimum area Santos:2021orr ; Santos:2023mee ; Santos:2024zoh . Thus, the “open wormhole” geometry Takayanagi:2011zk , if it exists, would be responsible for the entanglement of $S^{L}_{bdry}$ and $S^{R}_{bdry}$, both connected by a HM surface Hartman:2013qma (see Fig. 11).
Figure 11: Schematic form of Hartman-Maldacena (HM) surface connecting boundary field theory, which consists of two copies of a holographic CFT with BCFT boundary. In the right bipartition the time coordinate $t\to-t+i\beta/2$ was considered, which corresponds to a reversal of the time-like Killing vector field on the other side of the horizon of the black string.
4 Page curve behavior
The relation between entropy $S$ and area $A$ in the context of quantum gravity Hawking:1975vcx ; Santos:2021orr ; Sokoliuk:2022llp ; Santos:2023flb ; Santos:2023mee , particularly with a negative cosmological constant Maldacena:1997re ; Witten:1998qj , indeed connects deeply the holographic principle and the Ryu-Takayanagi conjecture Ryu:2006bv . Identifying the entanglement entropy with the area of minimal surfaces in the bulk serves as a bridge between gravitational theories and quantum information theory. This conjecture states that the entanglement entropy of a region in a quantum field theory is proportional to the area of a minimal surface in the corresponding bulk gravitational theory. For a black hole, this translates into the understanding that the entropy of the black hole is connected to the area of its event horizon. The consideration that the Hilbert space describing black hole dynamics is finite-dimensional, composed by orthogonal states often called “black hole microstates”, is significant Iliesiu:2022kny ; Balasubramanian:2022gmo . It suggests that despite the seemingly infinite degrees of freedom in a gravitational theory, the effective description can be captured within a finite framework, resembling the structure of statistical mechanics. Furthermore, the equivalence of quantum gravity with two AdS asymptotic boundaries, to two copies of BCFT, is particularly interesting. Each boundary (BCFTL and BCFTR) can be viewed as encoding information about the gravitational system, leading to a richer understanding of entanglement and correlation between the two sides. The semiclassical approach to calculate the entropy of black holes suggests that one can account for the information paradox through the evolution of the Page curve during black hole evaporation. This curve reflects how the entanglement entropy changes as the black hole looses mass, connecting quantum mechanics, thermodynamics, and information theory. Thus, the semiclassical gravitational path integral is not merely a computational tool but offers a convenient framework to explore the full Hilbert space dynamics. This perspective may help in understanding how information is preserved even when black holes evaporate Balasubramanian:2022gmo , addressing longstanding concerns about the fate of information in gravitational collapse.
The investigation of the Page curve in the context of Horndeski gravity, particularly through the AdS/BCFT correspondence, explores black hole information. The Page curve describes how a black hole entanglement entropy $S$ evolves during its evaporation. Initially, the entropy increases as the black hole absorbs matter, then it decreases as information is emitted. As a subclass of scalar-tensor theories, Horndeski gravity includes terms that account for higher-order derivatives of the scalar field. This framework allows for a richer set of dynamics than standard general relativity, making it relevant for understanding modifications to gravitational behavior in the presence of black holes.
The parameters $\alpha$ and $\gamma$ of Horndeski gravity affect the dynamics of the black hole and the resulting Page curve, altering spacetime geometry and affecting the behavior of the extremal surfaces, which contribute to entanglement entropy. Analyzing how these parameters affect the Page time (the moment at which the entropy begins to decrease) and Page angle (the steepness of the entropy curve during the evaporation process) is critical for understanding the interplay between black hole thermodynamics and quantum information Geng:2021mic ; Geng:2023qwm ; Geng:2021hlu . In summary, exploring the Page curve within Horndeski gravity offers a framework for understanding black hole information dynamics and contributes to the ongoing efforts to reconcile quantum mechanics with gravitational phenomena.
Figure 12: The Page angle $\rho_{Page}$ for various values of $\gamma$ and fixed $\alpha=-8/3$. The points where the Page angle gives the tensionless brane $\rho=0$ are those where the parameter $\gamma$ is furthest from zero. The fact that $\gamma$ is far from its null value shows that for $\rho_{Page}>0$ the island dominates at $t=0$, while for $\rho_{Page}<0$ the HM surface dominates at $t=0$. The regions showing dominant competition between the island and HM surfaces initially dominate the entropy calculation due to fixing $\alpha$ and varying $\gamma$, and reveal the strength of the scalar field contribution. These graphs are equivalent with the density plot along a diagonal slice in Fig. 14 below, with $r_{L}=r_{R}$.
[…]
To find the Page time, we need to impose that $S_{HM}(t_{Page})=S_{island}$, which provides
$t_{Page}=\frac{\beta}{4\pi}cosh^{-1}\left[\frac{e^{-2\rho_{*}}r_{L}r_{R}-r_{h}(\Delta_{L}+\Delta_{R})}{2\omega(\rho_{*})r_{h}\sqrt{\Delta_{L}\Delta_{R}}}\right],$ (74)
and considering the special case $r_{L}=r_{R}=r$ we have
$t_{Page}=\frac{\beta}{4\pi}cosh^{-1}\left[\frac{e^{-2\rho_{*}}r^{2}}{2\omega(\rho_{*})r_{h}\Delta_{r}}-\frac{1}{\omega(\rho_{*})}\right],$ (75)
$\Delta_{r}=r_{h}-r;\,\omega(\rho_{*})=1+\left(\frac{\alpha-3\gamma}{\gamma}\right)\cosh^{2}(\rho_{*}).$ (76)
Figure 13: The Page angle $\rho_{Page}$ for various values of $\alpha$ and fixed $\gamma=-0.4$. The points where the Page angle gives the tensionless brane $\rho=0$ are those where the parameter $\alpha$ is furthest from zero. The fact that $\alpha$ is far from its null value shows that for $\rho_{Page}>0$ the island dominates at $t=0$, while for $\rho_{Page}<0$ the HM surface dominates at $t=0$. The regions showing dominant competition between the island and HM surfaces initially dominate the entropy calculation due to fixing $\gamma$ and varying $\alpha$, and reveal the strength of the scalar field contribution. These graphs are equivalent with the density plot along a diagonal slice in Fig. 15 below, with $r_{L}=r_{R}$.
Now, in order to find the Page angle, we consider the condition $S_{HM}(t=0)=S_{island}$, which yields
$\rho_{Page}=\frac{1}{2}log\left[\frac{r_{L}r_{R}}{r_{h}\left(\Delta_{L}+\Delta_{R}+2\left[1+\left(\frac{\alpha-3\gamma}{\gamma}\right)\right]\sqrt{\Delta_{L}\Delta_{R}}\right)}\right],$ (77)
and for the special case $r_{L}=r_{R}=r$ we result to the expression
$\rho_{Page}=\frac{1}{2}log\left\{\frac{r^{2}}{r_{h}\left(2\Delta_{r}+2\left[1+\left(\frac{\alpha-3\gamma}{\gamma}\right)\right]\Delta_{r}\right)}\right\}.$ (78)
Figure 14: The evolution of the density of the Page angle $\rho_{Page}$ around $r_{L}$ and $r_{R}$, for various values of $\gamma$ and fixed $\alpha=-8/3$. The contour with $\rho_{Page}=0$ is depicted by a white dotted curve as a reference.
In Figs. 12 and 13 we depict the evolution of the Page angle density $\rho_{Page}$ as a function of $r_{L}/r_{h}$ and $r_{R}/r_{h}$, where $r_{L}$ and $r_{R}$ are the positions of the left and right boundaries, for various values of $\gamma$ and $\alpha$. These figures demonstrate that the Horndeski parameters can induce significant changes in the system, particularly when the radial coordinate $r$ is near 00 or the horizon radius $r_{h}$ Geng:2022dua . Such changes align with the understanding that Horndeski gravity causes a shift in the islands connected by the Hartman-Maldacena surface Hartman:2013qma . Our findings reveal that the residual information, denoted as $S^{L}_{bdry}$ and $S^{R}_{bdry}$, significantly affects the location of the bipartition, except when it is near the event horizon or a defect. Moreover, these figures illustrate the dependence of the Page curve behavior on the Horndeski parameters $\alpha$ and $\gamma$, as well as on the geometric factors $r_{L}$ and $r_{R}$. Finally, they provide visual evidence of the significant role that Horndeski gravity plays in modifying the dynamics of entanglement entropy compared to standard general relativity.
Figure 15: The evolution of the density of the Page angle $\rho_{Page}$ about $r_{L}$ and $r_{R}$, for various values of $\alpha$ and fixed $\gamma=-0.4$. The contour with $\rho_{Page}=0$ is depicted by a white dotted curve as a reference.
In Figs. 14 and 15 we present the evolution of the density of the Page angle $\rho_{Page}$ about $r_{L}$ and $r_{R}$, for various values of $\gamma$ and $\alpha$. The horizontal axis represents $r/r_{h}$, ranging from 0 to 1, the vertical axis shows the Page angle $\rho_{\star}$, and the color gradient indicates the logarithm of the Page time. The white dotted curves depict the contour where $\rho_{Page}=0$, which represents the transition point between island dominance ($\rho_{Page}>0$) and HM surface dominance ($\rho_{Page}<0$) at $t=0$. As $\gamma$ changes from $-1$ to $-0.1$, we observe significant changes in the density patterns, particularly near the corners of the plot where $r_{L}/r_{h}$ or $r_{R}/r_{h}$ approach 00 or $1$. These changes demonstrate the strong influence of the Horndeski parameter $\gamma$ on the Page angle, especially in regions close to the defect or the horizon. Regions where $\rho_{\star}>0$ indicate island dominance at $t=0$, while $\rho_{\star}<0$ shows Hartman-Maldacena surface dominance. As $\gamma$ approaches zero, we observe more pronounced variations in the Page angle, particularly for smaller values of $r/r_{h}$. Additionally, as $\alpha$ changes from $-11/3$ to $-1$, we observe significant changes in the density patterns, particularly in regions where $r_{L}/r_{h}$ or $r_{R}/r_{h}$ are close to 00 or $1$.
The above figures demonstrate that the Page time is highly sensitive to the bipartition point. This sensitivity can be understood as a consequence of the system geometry on the HM surface. Near the defect, as illustrated in Fig. 4, the lines are compressed by the influence of Horndeski scalar field, causing the metric to amplify the difference between the areas of the island and the HM surface. Conversely, when the HM surface approaches zero area (namely where both bipartitions simultaneously approach the event horizon) the surface extending from the event horizon to any boundary point retains a nonzero area. Consequently, the Page time experiences significant deviations when the degrees of freedom for small values of $\alpha$ and $\gamma$ in both bipartitions are limited, a scenario applicable only within the curved bottom theory.
5 Conclusions and discussion
In this work we investigated the interplay between Horndeski gravity, boundary conditions, and entanglement entropy, particularly within the Anti-de-Sitter/Boundary Conformal Field Theory (AdS/BCFT) correspondence framework. The modifications to the Page curve due to Horndeski gravity reveal the role of geometry in black hole information dynamics, potentially shedding light on the black hole information paradox. Moreover, the concept of entanglement islands, especially in the context of BCFT, can be helpful in examining quantum information distribution in these systems Andrzejewski:2023dja ; MohammadiMozaffar:2024uyp . As we discussed, the dependence of the Page time on Horndeski parameters can serve as a distinguishing feature for different gravity theories, opening the way for possible observational signatures.
Our analysis revealed that the Page curve depends on the Horndeski parameters $\alpha$ and $\gamma$. The fact that the Page time exhibits significant deviations when the number of degrees of freedom is small indicates the sensitivity of the entanglement structure to these parameters, particularly in curved backgrounds. This suggests that gravitational modifications can alter the fundamental aspects of quantum information dynamics. Our comparison with the case where the boundary is flat, where the HM surface area remains constant for symmetric bipartitions, reveals how the geometry affects the entanglement landscape, and in particular we saw that in curved geometries the interplay between the horizon and the boundary can lead to richer and more complex behaviors. Hence, investigating these dynamics in various geometries and parameter regimes could further elucidate the connections between gravity, quantum information, and holography.
In summary, we showed that Horndeski parameters significantly alter the behavior of the Page curve compared to standard general relativity, a feature caused by the nontrivial geometry induced by the scalar Horndeski field. Interestingly enough, the geometry far from the AdS limit plays a more significant role comparing to previous studies. This suggests that Horndeski gravity introduces important modifications to the distribution of quantum information in the holographic model. Lastly, we mention that the holographic consistency can be used reversely to impose constraints on Horndeski gravity itself, providing a new tool for probing the validity of modified gravity theories, establishing a novel connection between holography and the structure of viable gravitational theories.
Acknowledgements.
The authos would like to thank Hao Geng for the fruitful discussions. ENS acknowledges the contribution of the LISA CosWG, and of COST Actions CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)” and CA23130 “Bridging high and low energies in search of quantum gravity (BridgeQG)”.